OpenStudy (anonymous):

help with fractions: - 1/7 x (-7)/1=? not sure how to make this look right on here, but the first fraction is a negative one seventh, but the second fraction just shows the top number as a negative, not the bottom. so the second number reads as negative seven, positive oneths. I came up with 7/(-7) so is that a negative 1?

OpenStudy (phi):

Is this the problem \[\frac{-1}{7} \cdot \frac{-7}{1}\]

OpenStudy (phi):

To multiply fractions you multiply the tops -1 x -7 = 7, and then the bottoms 7*1= 7 the answer is 7/7 = 1

OpenStudy (anonymous):

it's showing the first fraction with the negative in front of the whole thing

OpenStudy (phi):

OK. These are all the same number. \[-\frac{7}{1}= \frac{-7}{1}= \frac{7}{-1}\]

OpenStudy (anonymous):

=1